I have a question that states: Find a function which is neither odd nor even.
a)x^4+9x+2+/x-2/
b)-11x^3+9x
c)5x^2-/x^2/+15
d)x^5+9x^3+5x
I cannot get this no matter what I try

Answer by stanbon(75887)

Find a function which is neither odd nor even.
--------
Comment:
f is even if f(x)=f(-x)
f is odd if f(x)=-f(-x)
For each of the problems you need to find f(-x) and maybe -f(-x) and
compare them to f(x)
----------
a)x^4+9x+2+/x-2/
f(-x) = (-x)^4 + 9(-x) + 2 + |-x-2|
f(-x) = x^4 -9x +2 + |-x-2|
--
-f(-x) = -x^4+9x-2-|-x-2|
It appears that neither equal f(x) so f is neither even nor odd.
---------------
b)-11x^3+9x
f(-x)=11x^3-9x
-f(-x)=-11x^3+9x
So, f is odd
-----------------------
c)5x^2-/x^2/+15
f(-x)=5x^2-|x^2|+15
So, f is even
---------------
d)x^5+9x^3+5x
f(-x)=-x^5-9x^3-5x
-f(-x)= x^5+9x^3+5x
So, f is odd
===================
Cheers,
Stan H.